OpenStudy (anonymous):

Calculus Integral 3 to -3 Square root 9-x^2 dx

OpenStudy (abb0t):

is this your integral? \[\int\limits_{-3}^{3} \sqrt{9-x^2}dx\]

OpenStudy (anonymous):

yes :o

OpenStudy (raden):

looks it is a circle with the centre is (0,0) and its radius 3

OpenStudy (raden):

so, we can solve without by integral

OpenStudy (raden):

just use the formula the area of circle : A = pi r^2 here r=3

OpenStudy (anonymous):

oh so you just look at the graph

OpenStudy (raden):

actually, yes :)

OpenStudy (anonymous):

how is the answer 9pi/2 though? where does the 2 come from

OpenStudy (abb0t):

It's not a circle by the way. It's a semi-circle.

OpenStudy (raden):

oh yeah, it just a semi circle :)

OpenStudy (raden):

btw, my assume is the curve y = sqrt(9-x^2) squaring to both sides : y^2 = 9-x^2 or x^2 + y^2 = 9 this, an equation of circle right ?

OpenStudy (anonymous):

ok, this makes more sense now, thank you guys :)